### v1

@@ Abstract
Convective processes are crucial in shaping exoplanetary atmospheres but are computationally expensive to simulate directly. A novel technique of simulating moist convection on tidally locked exoplanets is to use a global 3D model with a stretched mesh. This allows us to locally refine the model resolution to 4.7 km and resolve fine-scale convective processes without relying on parameterizations. We explore the impact of mesh stretching on the climate of a slowly rotating TRAPPIST-1e-like planet, assuming it is 1:1 tidally locked. In the stretched-mesh simulation with explicit convection, the climate is 5 K colder and 25% drier than that in the simulations with parameterized convection. This is due to the increased cloud reflectivity and exacerbated by the diminished greenhouse effect due to less water vapor. At the same time, our stretched-mesh simulations reproduce the key characteristics of the global climate of tidally locked rocky exoplanets, without any noticeable numerical artifacts. Our methodology opens an exciting and computationally feasible avenue for improving our understanding of 3D mixing in exoplanetary atmospheres. Our study also demonstrates the feasibility of a global stretched mesh configuration for LFRic-Atmosphere, the next-generation Met Office climate and weather model.
@@ 2 Methods
We use LFRic-Atmosphere, the new 3D GCM of the Met Office (Adams et al., 2019), based on a novel dynamical core GungHo and a suite of well-tested physical parameterizations inherited from its forerunner, the Unified Model (UM). GungHo solves the fully compressible non-hydrostatic Euler equations on a quasi-uniform cubed-sphere mesh using a mimetic finite-element discretisation and a mass-conserving finite-volume transport scheme (Melvin et al., 2019; Bendall et al., 2020, 2022; Kent et al., 2023). LFRic-Atmosphere reproduces a variety of atmospheric flows (Melvin et al., 2019; Kent et al., 2023; Brown et al., 2023; Melvin et al., 2024), including climate benchmarks for terrestrial exoplanets (Sergeev et al., 2023).
@@
Throughout this study we use the C192 cubed-sphere mesh, in a stretched and non-stretched configuration (see Appendix A for more details on mesh stretching). We performed four experiments: three with the mesh stretched by a factor of 10, and one with a non-stretched, quasi-uniform mesh (Fig 1), Appendix A). In the UniformParam and StretchParam configuration we use the same parameterizations as those used in Sergeev et al. (2023), including the standard mass-flux parameterization of convection. The StretchParam experiment differs from UniformParam only in the mesh configuration, while the StretchReduced experiment uses a reduced convective parameterization (Tomassini et al., 2023). The parameterization changes which are included in StretchReduced (for details, see Appendix B) are designed to allow small, unresolved convection to still be handled by the parameterization, whilst allowing larger and deeper convection to be explicitly resolved, and thus serves as an intermediate step between fully parameterized and fully explicit convection. The StretchReduced configuration, which can also be thought of as a “shallow-only” convection scheme, thus serves as an intermediate step between fully parameterized and explicit convection. The StretchExplicit configuration handles convection explicitly, i.e. with the parameterization switched off (see discussion in Sergeev et al., 2020).
@@ 3 Results
The stretched-mesh simulation without a convection parameterization (StretchExplicit), reproduces the so-called “cloud streets” on the day side’s periphery (Fig. 2a). These parallel elongated bands of clouds with cloud-free conditions between them were reported by Yang et al. (2023) in quasi-global convection-permitting simulations for TRAPPIST-1e. They form as a result of the persistent advection of cold air from the night side onto a warmer ocean surface of the day side, resulting in a formation of convective boundary layers. This is similar to the roll convection occurring in cold-air outbreaks on Earth (e.g. Gryschka & Raasch, 2005). Our work supports the results from Yang et al. (2023) and demonstrates that cloud streets can be reproduced in a stretched-mesh global model at a reduced computational cost.
@@
The small-scale variability enabled by the mesh refinement and explicit convection is also reflected in a marked increase in the precipitation rate in the substellar region. As Fig. 2h shows, the maximum instantaneous precipitation reaches in the UniformParam simulation. Moving from a quasi-uniform to a stretched mesh has little effect (Fig. 2g), while reducing or disabling the convection scheme results in almost doubling of the maximum precipitation rate (Fig. 2e,f). The histogram in Fig. 2i distills this change in the precipitation maxima. It shows that the mesh stretching barely changes precipitation rates (after interpolating the data to the same mesh). However, reducing the impact of the convection scheme or even entirely switching it off results in substantially more intense precipitation, reaching at least twice within our analysis period of 500 days ( of 221184 mesh cells). This is because in the StretchExplicit and StretchReduced cases much of the convection is handled by the grid-scale dynamics, permitting the formation of small but concentrated storms. Similarly high precipitation rate was reported for the stretched-mesh simulations of Earth as an aquaplanet Harris et al. (2016) and for the convection-permitting limited-area simulations of the climate of TRAPPIST-1e and Proxima Centauri b (Sergeev et al., 2020).
@@
The main effects of explicit convection on the global climate are a reduction in atmospheric water vapor content and an increase in condensed cloud water. While the UniformParam and StretchParam cases have a similar amount of water vapor (), in the StretchReduced and StretchExplicit cases it decreases by a about , with the latter simulation being the driest overall (Fig. 3a). The overall drying of the atmosphere in the StretchExplicit and StretchReduced simulations is evident in the vertical profiles of absolute humidity (Fig. 3b) and relative humidity (Fig. 3c). The relative humidity is noticeably lower in the troposphere in the StretchExplicit and StretchReduced simulations, so the water vapor decrease cannot be explained only by the lower temperatures in these two simulations (Fig. 3a). Indeed, since the stratospheric temperature is broadly the same in all four cases, it is the configuration of the convection scheme in the StretchReduced case that yields a less efficient upward transport of water vapor. This leads to a lower humidity in the upper layers of the night side and terminator regions (not shown), which may have a small effect on the water features in the transmission spectrum as discussed e.g. in Fauchez et al. (2022) and Sergeev et al. (2022a).
@@
The thermal maps in our simulations exhibits a day-night dichotomy typical for tidally locked planets, both in terms of surface temperature (Fig. 5a–d) and top-of-atmosphere outgoing longwave radiation (Fig. 5e–h). The substellar area reaches temperatures of , while the night-side coldest areas (stationary gyres) reach temperatures of . The latter substantially lower than the value predicted by a coarse-mesh LFRic-Atmosphere simulations of the THAI Hab 1 case (Sergeev et al., 2023), which is likely due to the slower rotation rate of the planet in the present study. Likewise, the outgoing longwave radiation is at the lower end of the THAI inter-model spread (Sergeev et al., 2022b): its values range from in the UniformParam experiment to in StretchExplicit (Fig. 5e–h). Clouds prevent even more heat being lost to space, so their net effect on the top-of-atmosphere longwave radiation flux is positive. It is broadly the same in our four experiments () and is comparable to that predicted by the models in Sergeev et al. (2022b) and Yang et al. (2023).
@@ 4 Discussion
Our results show that LFRic-Atmosphere is capable of simulating small-scale atmospheric processes using a stretched mesh while also reproducing the key characteristics of the global climate of a tidally locked terrestrial exoplanet. Using a locally refined mesh allows us to capture small-scale cloud patterns in the substellar region. This includes the banded cloud structures on the day side of the planet, known as “cloud streets” previously examined in convection-permitting simulations in Yang et al. (2023). In agreement with previous studies based on localized or global high-resolution grids (e.g. Harris et al., 2016; Kajikawa et al., 2016; Uchida et al., 2016; Rios‐Berrios et al., 2022), resolving deep convection leads to more intense and localized precipitation. Thus, at a reduced computational cost (Table 1), LFRic-Atmosphere allows us to examine a regional phenomenon such as convection in great detail and at the same time study the interaction between local and global atmospheric circulation. Tidally locked exoplanets offer a particularly suitable use case but a locally refined mesh has also shown potential for Mars climate modeling, e.g. at the scale of individual craters (Lian & Richardson, 2023).
@@
Our study is the first exoplanet application of LFRic-Atmosphere in a global configuration with explicit convection, allowed for by the stretched mesh. Our key finding is that the simulation with explicit convection (StretchExplicit) predicts a climate that is substantially drier (as seen in both specific and relative humidity) and colder (by ), with about more cloud condensate compared to the simulations with fully parameterized convection. The main reason for the colder climate is the higher cloud albedo, i.e. stronger reflection of the shortwave radiation by clouds (by ). Additionally, there is less water vapor and thus a weaker greenhouse effect. The increase in cloud condensate, especially in the low-level cloud liquid, can be partly explained by the fact that the grid resolution on the day side’s periphery may be too coarse to fully resolve the shallow convection. However, the experiment with reduced parameterization (StretchReduced) produces a similar cloud structure (between that in the StretchExplicit run and those with fully parameterized convection), supporting this trend. We believe that overall our experiment with explicit convection produces a simulated climate closer to reality because it truly represents the underlying physics of the atmosphere, without relying on parameterizations tuned for Earth. In the absence of observational constraints, these simulations may serve as benchmarks for coarse-resolution GCMs with parameterized convection. At the same time, as noted in Tomassini et al. (2023), cloud parameterizations will likely require tuning for convection-permitting models. This may reduce the amount of shortwave radiation reflected by the clouds, resulting in a warmer climate. Future work should explore the representation of cloud microphysics and their impact on the climate in stretched-mesh simulations.
@@
A caveat of the present study is that the convection scheme in our model is not tuned for a stretched mesh configuration. Indeed, in most km-scale models convection is only partially resolved (Tomassini et al., 2023). This introduces the challenge of representing convection in a balanced way: partially explicitly and partially as a parameterized sub-grid process. Scale-aware convection schemes are a promising solution (Kendon et al., 2021), which may also re-ignite interest in using stretched-mesh GCMs (Uchida et al., 2016). One example of a scale-aware scheme is CoMorph, a new convection parametrization under development at the Met Office (Daleu et al., 2023; Lavender et al., 2024). It is designed for use within both the Unified Model and LFRic-Atmosphere and has already shown to be close to cloud-resolving models in representing many aspects of convection (Lavender et al., 2024). The benefits of applying CoMorph in a stretched-mesh configuration will be explored for a tidally locked exoplanet in a future study.
@@ Appendix B Convection and turbulence parameters in the StretchReduced experiment
In the StretchReduced experiment, we follow a km-scale configuration used for the Unified Model in Tomassini et al. (2023), namely their MidLevShConv15RAturb configuration. This configuration combines a reduced convection parameterization with a modified representation boundary layer scheme. As a result, convection is allowed to be partly explicit, accounting for the fact that in the focal region of our stretched mesh the grid step reaches .
Briefly, the StretchReduced parameterization settings are different to those used in StretchParam and UniformParam in the following. Scaling of the shallow convective mass flux is set to 0.015 (default: 0.03) — the value found to produce more realistic low-level clouds in global km-scale numerical weather prediction runs. The formulation of the mid-level convection scheme is also modified. The key modification that reduces the sub-grid convective mass flux according to the model resolution, allowing convection to be partly explicit, is a longer timescale for the convective available potential energy (CAPE) closure ( instead of the default ). The configuration also includes the turbulence blending scheme (Boutle et al., 2014), which allows for a transition between the 1D turbulence parameterization (used in coarse-grid GCMs) with a 3D Smagorinsky-Lilly turbulence scheme (used in cloud-resolving models).

### v2

@@ Abstract
Convective processes are crucial in shaping exoplanetary atmospheres but are computationally expensive to simulate directly. A novel technique of simulating moist convection on tidally locked exoplanets is to use a global 3D model with a stretched mesh. This allows us to locally refine the model resolution to 4.7 km and resolve fine-scale convective processes without relying on parameterizations. We explore the impact of mesh stretching on the climate of a slowly rotating TRAPPIST-1e-like planet, assuming it is 1:1 tidally locked. In the stretched-mesh simulation with explicit convection, the climate is 5 K colder and 25% drier than that in the simulations with parameterized convection (with both stretched and quasi-uniform meshes). This is due to the increased cloud reflectivity — because of an increase of low-level cloudiness — and exacerbated by the diminished greenhouse effect due to less water vapor. At the same time, our stretched-mesh simulations reproduce the key characteristics of the global climate of tidally locked rocky exoplanets, without any noticeable numerical artifacts. Our methodology opens an exciting and computationally feasible avenue for improving our understanding of 3D mixing in exoplanetary atmospheres. Our study also demonstrates the feasibility of a global stretched mesh configuration for LFRic-Atmosphere, the next-generation Met Office climate and weather model.
@@ 2 Methods
We use LFRic-Atmosphere, the new 3D GCM of the Met Office (Adams et al., 2019), based on a novel dynamical core GungHo and a suite of well-tested physical parameterizations inherited from its forerunner, the Unified Model (UM). GungHo solves the fully compressible non-hydrostatic Euler equations on a quasi-uniform cubed-sphere mesh using a mimetic finite-element discretisation and a mass-conserving finite-volume transport scheme (Melvin et al., 2019; Bendall et al., 2020, 2022; Kent et al., 2023). The radiative transfer is parameterized using a two-stream correlated- scheme SOCRATES based on Edwards & Slingo (1996), for which we use the same setup as in Sergeev et al. (2022b). LFRic-Atmosphere reproduces a variety of atmospheric flows (Melvin et al., 2019; Kent et al., 2023; Brown et al., 2023; Melvin et al., 2024), including climate benchmarks for terrestrial exoplanets (Sergeev et al., 2023).
@@
Throughout this study we use the C192 cubed-sphere mesh, in a stretched and non-stretched configuration (see Appendix A for more details on mesh stretching). We performed four experiments: three with the mesh stretched by a factor of 10, and one with a non-stretched, quasi-uniform mesh (Fig 1), Appendix A). In the UniformParam and StretchParam configuration we use the same parameterizations as those used in Sergeev et al. (2023), including the standard mass-flux parameterization of convection. The StretchParam experiment differs from UniformParam only in the mesh configuration, while the StretchReduced experiment uses a “reduced” convective parameterization (for technical details, see Appendix B). The StretchReduced configuration is designed to allow small, unresolved convection to be handled by the parameterization, whilst allowing deeper convection to be explicitly resolved. Note that the transition between resolved and parameterized convection in StretchReduced is not defined by a specific threshold. Instead, it dynamically changes with the cell size (and is therefore particularly useful in a stretched-mesh setup): when the scale of convection is smaller than the grid scale, convection is handled by the parameterization, whilst when the scale of convection is greater than the grid scale, parameterization effectively switches off and convection is handled explicitly. The StretchReduced experiment thus serves as an intermediate step between fully parameterized and fully explicit convection. The StretchExplicit configuration handles convection explicitly, i.e. with the parameterization switched off (see discussion in Sergeev et al., 2020).
@@ 3 Results
On the periphery of the substellar region, the stretched-mesh simulation without a convection parameterization (StretchExplicit), produces cloud bands reminiscent of “cloud streets” (Fig. 2a). These parallel elongated bands of clouds with cloud-free conditions between them were reported by Yang et al. (2023) in quasi-global convection-permitting simulations for TRAPPIST-1e. They form as a result of the persistent advection of cold air from the night side onto a warmer ocean surface of the day side, resulting in a formation of convective boundary layers. This is similar to the roll convection occurring in cold-air outbreaks on Earth (e.g. Gryschka & Raasch, 2005). Cloud streets can have a cooling effect on the global climate because the outgoing longwave radiation is larger in the clear-sky regions than that in the cloudy regions (Yang et al., 2023). Our work supports the results from Yang et al. (2023) and suggests that cloud streets can be reproduced in a stretched-mesh global model at a reduced computational cost. However, they may also be an artifact of under-resolving turbulence because of the increasing grid spacing and the 1D boundary layer parameterization in our setup (for the relevant discussion, see Boutle et al., 2014).
@@
The small-scale variability enabled by the mesh refinement and explicit convection is also reflected in a marked increase in the precipitation rate in the substellar region. As Fig. 2h shows, the maximum instantaneous precipitation reaches in the UniformParam simulation. Moving from a quasi-uniform to a stretched mesh has little effect (Fig. 2g), while reducing or disabling the convection scheme results in almost doubling of the maximum precipitation rate (Fig. 2e,f). The histogram in Fig. 2i distills this change in the precipitation maxima. It shows that the mesh stretching barely changes precipitation rates (after interpolating the data to the same mesh). However, reducing the impact of the convection scheme or even entirely switching it off results in substantially more intense precipitation, reaching at least twice within our analysis period of 500 days ( of 221184 mesh cells). This is because in the StretchExplicit and StretchReduced cases much of the convection is handled by the grid-scale dynamics, permitting the formation of small but concentrated storms. Similarly high precipitation rate was reported for the stretched-mesh simulations of Earth as an aquaplanet (Harris et al., 2016) and for the convection-permitting limited-area simulations of the climate of TRAPPIST-1e and Proxima Centauri b (Sergeev et al., 2020).
@@
The main effects of explicit convection on the global climate are a reduction in atmospheric water vapor content and an increase in condensed cloud water. While the UniformParam and StretchParam cases have a similar amount of water vapor (), in the StretchReduced and StretchExplicit cases it decreases by a about , with the latter simulation being the driest overall (Fig. 3a). The overall drying of the atmosphere in the StretchExplicit and StretchReduced simulations is evident in the vertical profiles of absolute humidity (Fig. 4b) and relative humidity (Fig. 4c). The relative humidity is noticeably lower in the troposphere in the StretchExplicit and StretchReduced simulations, so the water vapor decrease cannot be explained only by the lower temperatures in these two simulations (Fig. 4a). Indeed, since the stratospheric temperature is broadly the same in all four cases, it is the parameterization of turbulent mixing above the boundary layer included in the StretchReduced configuration that reduces the mixing across the tropopause and results in a less efficient upward transport of water vapor. This leads to a lower humidity in the upper layers of the night side and terminator regions (not shown), which may have a small effect on the water features in the transmission spectrum as discussed e.g. in Fauchez et al. (2022) and Sergeev et al. (2022a).
@@
Compared to previous high-resolution studies of e.g. Sergeev et al. (2020), our results broadly agree that the high-resolution model with explicit convection produces more low-level clouds than a coarse-resolution GCM with parameterized convection. This is unsurprising because most subgrid-scale physics parameterizations in LFRic-Atmosphere are inherited from the Unified Model (Sergeev et al., 2023). The differences between our study and that of Lefèvre et al. (2021) are greater mostly because of the inter-model differences in physical parameterizations. Moreover, the high-resolution simulation in Lefèvre et al. (2021) is forced by a constant and spatially uniform heating rate, while in the present study, the impact of the global circulation on the substellar region is direct due to the gradual stretching of the mesh and no boundaries between nested and driving models. Finally, the planetary parameters are different between the aforementioned studies and the present study. As a result, the substellar atmosphere is warmer in the simulations in Lefèvre et al. (2021) and Sergeev et al. (2020) than in the present study. This likely leads to more water being in the vapor form than in the condensed cloud form in those two studies. Another consequence of a colder climate in our simulations is that the substellar convection is likely to be slightly weaker, resulting in a weaker transport of cloud condensate in the upper atmosphere and hence fewer high clouds.
@@
The thermal maps in our simulations exhibits a day-night dichotomy typical for tidally locked planets, both in terms of surface temperature (Fig. 5a–d) and top-of-atmosphere outgoing longwave radiation (Fig. 5e–h). The substellar area reaches temperatures of , while the night-side coldest areas (stationary gyres) reach temperatures of . The latter substantially lower than the value predicted by a coarse-mesh LFRic-Atmosphere simulations of the THAI Hab 1 case (Sergeev et al., 2023), which is likely due to the slower rotation rate of the planet in the present study. A similar effect of slowing the rotation rate on the climate was found in GCM simulations by e.g. Edson et al. (2011) and Komacek & Abbot (2019). Likewise, the outgoing longwave radiation is at the lower end of the THAI inter-model spread (Sergeev et al., 2022b): its values range from in the UniformParam experiment to in StretchExplicit (Fig. 5e–h). Clouds prevent even more heat being lost to space, so their net effect on the top-of-atmosphere longwave radiation flux is positive. It is broadly the same in our four experiments () and is comparable to that predicted by the models in Sergeev et al. (2022b) and Yang et al. (2023).
@@ 4 Discussion
Our results show that LFRic-Atmosphere is capable of simulating small-scale atmospheric processes using a stretched mesh while also reproducing the key characteristics of the global climate of a tidally locked terrestrial exoplanet. Using a locally refined mesh allows us to capture small-scale cloud patterns in the substellar region. In agreement with previous studies based on localized or global high-resolution grids (e.g. Harris et al., 2016; Kajikawa et al., 2016; Uchida et al., 2016; Rios‐Berrios et al., 2022), resolving deep convection leads to more intense and localized precipitation. Thus, at a reduced computational cost (Table 1), LFRic-Atmosphere allows us to examine a regional phenomenon such as convection in great detail and at the same time study the interaction between local and global atmospheric circulation.
Tidally locked exoplanets offer a particularly suitable use case but a locally refined mesh has also shown potential for Mars climate modeling, e.g. at the scale of individual craters (Lian & Richardson, 2023). Stretched-mesh GCMs may also be useful in studying Titan weather, in particular the interaction between localized convective storms and large-scale Rossby waves (Battalio & Lora, 2021). For Venus and slowly-rotating Venus-like exoplanets, stretched-mesh GCMs may shed light on the convectively-generated gravity waves and their interaction with the superrotating flow (Lefèvre et al., 2018). For non tidally locked exoplanets, the mesh can be stretched in a different way, e.g. focusing the high resolution on the equatorial band, allowing for a better simulation of deep convection in the tropics (Rios‐Berrios et al., 2022).
@@
Our study is the first exoplanet application of LFRic-Atmosphere in a global configuration with explicit convection, allowed for by the stretched mesh. Our key finding is that the simulation with explicit convection (StretchExplicit) predicts a climate that is substantially drier (as seen in both specific and relative humidity) and colder (by ), with about more cloud condensate compared to the simulations with fully parameterized convection. The main reason for the colder climate is the higher cloud albedo, i.e. stronger reflection of the shortwave radiation by clouds (by ). Additionally, there is less water vapor and thus a weaker greenhouse effect. The increase in cloud condensate, especially in the low-level cloud liquid, can be partly explained by the fact that the grid resolution on the day side’s periphery may be too coarse to fully resolve the shallow convection. However, the experiment with reduced parameterization (StretchReduced) produces a similar cloud structure (between that in the StretchExplicit run and those with fully parameterized convection), supporting this trend. Note that the StretchReduced simulation serves as an intermediate solution between the parameterized and explicit convection because it is designed to activate convection depending not only on the grid cell size but the size of convection relative to it.
We believe that overall our experiment with explicit convection simulates a climate close to “reality” because it truly represents the underlying physics of the atmosphere, without relying on parameterizations tuned for Earth. The main caveat is the grid coarsening on the day-side periphery where convection becomes under-resolved. A possible solution could be to use a base mesh with more grid points and a smaller stretching factor. In the absence of observational constraints, these simulations may serve as benchmarks for coarse-resolution GCMs with parameterized convection. At the same time, as noted in Tomassini et al. (2023), cloud parameterizations will likely require tuning for convection-permitting models. This may reduce the amount of shortwave radiation reflected by the clouds, resulting in a warmer climate. Future work should explore the representation of cloud microphysics and their impact on the climate in stretched-mesh simulations.
@@
Another caveat of the present study is that the convection scheme in our model is not tuned for a stretched mesh configuration. Indeed, in most km-scale models convection is only partially resolved (Tomassini et al., 2023). This introduces the challenge of representing convection in a balanced way: partially explicitly and partially as a parameterized sub-grid process. Scale-aware convection schemes are a promising solution (Kendon et al., 2021), which may also re-ignite interest in using stretched-mesh GCMs (Uchida et al., 2016). One example of a scale-aware scheme is CoMorph, a new convection parametrization under development at the Met Office (Daleu et al., 2023; Lavender et al., 2024). It is designed for use within both the Unified Model and LFRic-Atmosphere and has already shown to be close to cloud-resolving models in representing many aspects of convection (Lavender et al., 2024). The benefits of applying CoMorph in a stretched-mesh configuration will be explored for a tidally locked exoplanet in a future study.
@@ Appendix B Convection and turbulence parameters in the StretchReduced experiment
In the StretchReduced experiment, we follow a km-scale configuration used for the Unified Model in Tomassini et al. (2023), namely their MidLevShConv15RAturb configuration. This configuration combines a “reduced” convection parameterization with a modified representation boundary layer scheme. As a result, convection is allowed to be partly explicit, accounting for the fact that in the focal region of our stretched mesh the grid step reaches .
Briefly, the StretchReduced parameterization settings are different to those used in StretchParam and UniformParam in the following. Scaling of the shallow convective mass flux is set to 0.015 (default: 0.03) — the value found to produce more realistic low-level clouds in global km-scale numerical weather prediction runs. The formulation of the mid-level convection scheme is also modified. The key modification that reduces the sub-grid convective mass flux is a longer timescale for the convective available potential energy (CAPE) closure: instead of the default . This effectively decreases the strength of the convection parameterization and allows convection to be partly explicit, i.e. represented by grid-scale vertical flow. The configuration also includes the turbulence blending scheme (Boutle et al., 2014), which allows for a transition between the 1D turbulence parameterization (used in coarse-grid GCMs) with a 3D Smagorinsky-Lilly turbulence scheme (used in cloud-resolving models). As mentioned in Sec. 3, this parameterization leads to a drier stratosphere compared to the other three simulations.